# All-optical mapping of ultrafast carrier dynamics in a Dirac semimetal

### v1

@@ Conclusion
In conclusion, our joint experimental and theoretical study has established an HHG-based all-optical approach to observe and quantify ultrafast depletion effects and carrier excitation dynamics in solids, which also yields absolute timing information. With respect to conventional bulk materials, such as ZnO, we found that the Dirac semimetal HOPG shows strong valence band population saturation due to its gapless nature even under moderate laser intensities (), leading to half-filling in the vicinity of the Dirac point. We identified this effect as responsible for the observed fs time shifts and the suppression of the harmonic emission in the later part of the driving laser pulse through Pauli blocking.
Our study demonstrates that gapless materials are subject to depletion effects and Pauli blocking, leaving behind a sizable conduction band population and rendering harmonic generation processes nonparametric in nature. Petahertz optoelectronics, where optical manipulation of the material is performed on attosecond time scales, relies mainly on virtual charge carrier excitations and their reversibility 33, 34, 8. In order to realize petahertz optoelectronics with full reversibility in gapless materials, pulse duration and peak intensity need to be limited such that Pauli blocking is avoided. Even if this cannot be achieved, gapless materials can nevertheless serve as an ultrafast optoelectronic switch where the presence or the absence of a pre-excitation pulse affects the efficiency of nonlinear processes through Pauli blocking on the significantly longer time scale of spontaneous emission. The nonparametric nature of HHG in gapless materials even at moderate intensities may also be helpful for generating upconverted light with nonclassical photon statistics due to the depletion effect, as recently predicted for gas-phase HHG 35.
@@ Methods
In the experiment, the fundamental () pulse at 1980 nm is generated from a phase-stable optical parametric chirped pulse amplification (OPCPA) laser system operating at a repetition rate of 1 MHz (Class 5 Photonics White Dwarf 36). The pulse duration of the field measured by frequency-resolved optical gating (FROG) yields a full-width-at-half-maximum (FWHM) duration of approximately 60 fs ( optical cycles). The second harmonic () pulse was generated by focusing the fundamental beam into a -barium borate (BBO) crystal under optimum phase-matching conditions. By adjusting the power, we maintained the intensity ratio between the and pulses at . For HHG in HOPG, the pulse energy was set between 1–5 nJ, while for ZnO, a higher energy of approximately 30 nJ was used.
The two-color interferometer setup was implemented by first separating the and beams using a dichroic beam splitter. The relative delay between the two pulses was controlled via a piezo delay stage placed in the beam path. The two beams were then recombined by reflecting them off two mirrors back into the same dichroic beam splitter, with a slight lateral offset angle relative to the incoming paths, and subsequently directed toward the sample for HHG. A parabolic mirror with mm focuses the light onto the HOPG sample (TipsNano) at an angle of incidence of 14 degrees. The beam size at the focus was measured using a knife-edge method, yielding an effective radius of (see discussion and Fig. S1 in the Supplementary Information). We collect the harmonic light with a lens and send it to a Si-based spectrometer.
@@
To describe the interaction of the two-color field and HOPG in the Dirac cone, we use the following one-dimensional semiconductor Bloch equations37, 38 (SBEs) centered around the Dirac cone:
Here, represents a dimensionless polarization depending on time and momentum , induced by the time-dependent electric field . The indices and correspond to the valence band and conduction band, respectively. denotes the population of electrons (holes) in the conduction (valence) band. Here, we consider linear energy-momentum dispersion relation around Dirac cone with Fermi velocity describing a continuum ground state Dirac Hamiltonian. This approximation should be valid in our regime where the laser-driven electrons only move near the Dirac cone. The transition dipole moments (TDMs) are calculated as 39 (see Fig. S3 in the Supplementary Information). We set the phenomenological interband polarization decoherence time to one optical cycle of the 2 m driving laser field fs, and the relaxation time for intraband current to . This consideration is valid since the relaxation is generally long and ranges from few hundred femtosecond to picosecond in graphene and is less sensitive to the results 40, 41. Band structure and TDM for ZnO crystal along the –M direction are shown in Fig. S4 of the Supplementary Information (for details see the Supplementary Information). The calculations are performed with dense -grid of size 200, using a time step of 2.4 as and according to the experimental parameters. The yield of the harmonics for the two-color case are calculated with two-color delay steps of 0.33 fs.
The HHG spectrum is computed as 37
@@
denote the interband and intraband contributions to the HHG yield, respectively. represents the group velocity of the electrons and holes in the corresponding bands.
@@
For a theoretical analysis of saturation and electron population dynamics, we begin by considering the non-depleted excitation rate of electrons from the valence band to the conduction band. According to Fermi’s golden rule, this excitation rate at a given crystal momentum and time can be expressed as
The total excitation rate integrated over the Brillouin zone is then given by
@@
For interband-driven high-harmonic generation (HHG), the emitted harmonic intensity is proportional to the instantaneous excitation rate, .
@@
M. K. and M. F. C. conceived and supervised the project, joined by O. N. in the later stage of the project. Z. C. built the experimental setup and performed the measurements. C. G. performed initial simulations. Z. C. carried out the theory calculations and C. G. and O. N. performed independent supplemental calculations. Z. C. and C. G. wrote the initial manuscript. All authors contributed to the interpretation of the results and to the preparation of the final manuscript.
@@
Correspondence to Michael Krüger or Marcelo F. Ciappina.
@@ Acknowledgements.
We thank M. Ivanov, M. Segev, Q. Yan and Y. Wu for insightful discussions and U. Leonhardt and L. M. Procopio for providing specialized equipment. Z. C., C. G., I. N., D. K., M. F. C and M. K. acknowledge the Guangdong Technion – Israel Institute of Technology (GTIIT) and Technion Seed Grant Program for enabling their collaboration and joint research. Z. C., I. N., D. K., M. F. C and M. K. thank the Helen Diller Quantum Center at the Technion for partial financial support. M. F. C. acknowledges support by the National Key Research and Development Program of China (Grant No. 2023YFA1407100), Guangdong Province Science and Technology Major Project (Future functional materials under extreme conditions - 2021B0301030005) and the Guangdong Natural Science Foundation (General Program project No. 2023A1515010871).Ethics declarations

### v2

@@ Conclusion
In conclusion, our joint experimental and theoretical study has established an NPHG-based all-optical approach to observe and quantify ultrafast carrier saturation effects and carrier excitation dynamics in solids, which also yields absolute timing information. With respect to conventional bulk materials, such as ZnO, we found that the Dirac semimetal HOPG shows strong conduction band population saturation due to its gapless nature even under moderate laser intensities (), leading to half-filling in the vicinity of the Dirac point. We identified this effect as responsible for the observed fs time shifts and the suppression of the harmonic emission in the later part of the driving laser pulse through state blocking.
Our study demonstrates that gapless materials are subject to saturation effects and state blocking, leaving behind a sizable conduction band population and rendering harmonic generation processes nonparametric in nature. Petahertz optoelectronics, where optical manipulation of the material is performed on attosecond time scales, relies mainly on virtual charge carrier excitations and their reversibility 49, 50, 8. In order to realize petahertz optoelectronics with full reversibility in gapless materials, pulse duration and peak intensity need to be limited such that state blocking is avoided. Even if this cannot be achieved, gapless materials can nevertheless serve as an ultrafast optoelectronic switch where the presence or the absence of a pre-excitation pulse affects the efficiency of nonlinear processes through state blocking on the significantly longer time scale of spontaneous emission. The nonparametric nature of HHG in gapless materials even at moderate intensities may also be helpful for generating upconverted light with nonclassical photon statistics due to a saturation or depletion effect, as recently predicted for gas-phase HHG 51.
@@ Methods
In the experiment, the fundamental () pulse at 1980 nm is generated from a phase-stable optical parametric chirped pulse amplification (OPCPA) laser system operating at a repetition rate of 1 MHz (Class 5 Photonics White Dwarf 52). The pulse duration of the field measured by frequency-resolved optical gating (FROG) yields a full-width-at-half-maximum (FWHM) duration of approximately 60 fs ( optical cycles). The second harmonic () pulse was generated by focusing the fundamental beam into a -barium borate (BBO) crystal under optimum phase-matching conditions. By adjusting the power, we maintained the intensity ratio between the and pulses at . For HHG in HOPG, the pulse energy was set between 1–5 nJ, while for ZnO, a higher energy of approximately 30 nJ was used.
The two-color interferometer setup was implemented by first separating the and beams using a dichroic beam splitter. The relative delay between the two pulses was controlled via a piezo delay stage placed in the beam path. The two beams were then recombined by reflecting them off two mirrors back into the same dichroic beam splitter, with a slight lateral offset angle relative to the incoming paths, and subsequently directed toward the sample for harmonic generation. A parabolic mirror with mm focuses the light onto the HOPG sample (TipsNano) at an angle of incidence of 14 degrees. The beam size at the focus was measured using a knife-edge method, yielding an effective radius of (see discussion and Fig. S1 in the Supplementary Information). We collect the harmonic light with a lens and send it to a Si-based spectrometer.
@@
To describe the interaction of the two-color field and HOPG in the Dirac cone, we use the following one-dimensional semiconductor Bloch equations53, 54 (SBEs) centered around the Dirac cone:
Here, represents a dimensionless polarization depending on time and momentum , induced by the time-dependent electric field . The indices and correspond to the valence band and conduction band, respectively. denotes the population of electrons (holes) in the conduction (valence) band. Here, we consider linear energy-momentum dispersion relation around Dirac cone with Fermi velocity describing a continuum ground state Dirac Hamiltonian. This approximation should be valid in our regime where the laser-driven electrons only move near the Dirac cone. The transition dipole moments (TDMs) are calculated as 46 (see Fig. S3 in the Supplementary Information). We set the phenomenological interband polarization decoherence time to one optical cycle of the 2 m driving laser field fs, and the relaxation time for intraband current to . This consideration is valid since the relaxation is generally long and ranges from few hundred femtosecond to picosecond in graphene and is less sensitive to the results 47, 39. Band structure and TDM for ZnO crystal along the –M direction are shown in Fig. S4 of the Supplementary Information (for details see the Supplementary Information). The calculations are performed with dense -grid of size 200, using a time step of 2.4 as and according to the experimental parameters. The yield of the harmonics for the two-color case are calculated with two-color delay steps of 0.33 fs.
The HHG spectrum is computed as 53
@@
denote the interband and intraband contributions to the harmonic yield, respectively. represents the group velocity of the electrons and holes in the corresponding bands.
@@
For a theoretical analysis of saturation and electron population dynamics, we begin by considering the non-depleted excitation rate of electrons from the valence band to the conduction band. The excitation rate at a given crystal momentum and time can be expressed as
w_0(k, t) ∼2πℏ — ⟨ψ_c — E(t) ⋅^d — ψ_v ⟩—^2 ≈2π— E(t) ⋅d_cv(k+A(t)) —^2.
The total excitation rate integrated over the Brillouin zone is then given by {align} w(t) = ∫w_0(k, t) dk = 2πE^2(t) ∫—d_cv(k+A(t))—^2 dk.
@@
For interband-driven harmonic generation, the emitted harmonic intensity is proportional to the instantaneous excitation rate, .
@@
M. K. and M. F. C. conceived and supervised the project, joined by O. N. in the later stage of the project. Z. C. built the experimental setup and performed the measurements. C. G. performed initial simulations. Z. C. carried out the theory calculations and C. G. and E. U. and O. N. performed independent supplemental calculations. Z. C. wrote the initial manuscript. All authors contributed to the interpretation of the results and to the preparation of the final manuscript.
@@
Correspondence to Michael Krüger or Marcelo F. Ciappina or Ofer Neufeld.
@@ Acknowledgements.
We thank M. Ivanov, M. Segev, Q. Yan and Y. Wu for insightful discussions, U. Leonhardt and L. M. Procopio for providing specialized equipment, and X. Duan for providing the monolayer WS2 sample. Z. C., C. G., I. N., D. K., M. F. C. and M. K. acknowledge the Guangdong Technion – Israel Institute of Technology (GTIIT) and Technion Seed Grant Program for enabling their collaboration and joint research. Z. C., I. N., D. K., E. U., O. N. and M. K. thank the Helen Diller Quantum Center and the Russell Berrie Nanotechnology Institute at the Technion for partial financial support. M. F. C. acknowledges support by the National Key Research and Development Program of China (Grant No. 2023YFA1407100), Guangdong Province Science and Technology Major Project (Future functional materials under extreme conditions - 2021B0301030005), the Guangdong Natural Science Foundation (General Program project No. 2023A1515010871) and the National Natural Science Foundation of China (Grant No. 12574092).Ethics declarations